OpenStudy (anonymous):

Find the distance between points P(4, 6) and Q(8, 9) to the nearest tenth

OpenStudy (ivettef365):

use distance formula

OpenStudy (anonymous):

the website is invalid

OpenStudy (ivettef365):

it let me go to it, but try this one http://www.mathwarehouse.com/calculators/distance-formula-calculator.php

OpenStudy (anonymous):

here you go http://prntscr.com/1bwkox

OpenStudy (anonymous):

if you know the pythogorean theorem you can easily remember the distance formula since it comes from it. a^2+b^2=c^2 :) easier to remember.

OpenStudy (anonymous):

just sqrt bot sides c=sqrt(a^2+b^2) where a and b are your differences in x and y